\begin{thebibliography}{99}

\bibitem{MZ}
J.~A.~Minahan and K.~Zarembo,
``The Bethe-ansatz for N = 4 super Yang-Mills,''
JHEP {\bf 0303} (2003) 013
[arXiv:hep-th/0212208].

\bibitem{revs}
N.~Beisert,
``The dilatation operator of N = 4 super Yang-Mills theory and
integrability,''
Phys.\ Rept.\  {\bf 405} (2005) 1
[arXiv:hep-th/0407277].
$\bullet$
J.~Plefka,
``Spinning strings and integrable spin chains in the AdS/CFT
correspondence,''
arXiv:hep-th/0507136.
$\bullet$
J.~A.~Minahan,
``A Brief Introduction To The Bethe Ansatz In N=4 Super-Yang-Mills,''
J.\ Phys.\ A  {\bf 39}, 12657 (2006).
%%CITATION = JPAGB,A39,12657;%%

\bibitem{BeiSmat}
N.~Beisert,
``The su(2$|$2) dynamic S-matrix,''
arXiv:hep-th/0511082.

\bibitem{Jan}
R.~A.~Janik,
``The AdS(5) x S**5 superstring worldsheet S-matrix and crossing
symmetry,''
Phys.\ Rev.\ D {\bf 73} (2006) 086006
[arXiv:hep-th/0603038].

\bibitem{hopf}
C.~Gomez and R.~Hernandez,
``The magnon kinematics of the AdS/CFT correspondence,''
JHEP {\bf 0611} (2006) 021
[arXiv:hep-th/0608029].

\bibitem{PST}
J.~Plefka, F.~Spill and A.~Torrielli,
``On the Hopf algebra structure of the AdS/CFT S-matrix,''
Phys.\ Rev.\ D {\bf 74} (2006) 066008
[arXiv:hep-th/0608038].

\bibitem{pert}
N.~Beisert, C.~Kristjansen and M.~Staudacher,
``The dilatation operator of N = 4 super Yang-Mills theory,''
Nucl.\ Phys.\ B {\bf 664} (2003) 131
[arXiv:hep-th/0303060].
$\bullet$
N.~Beisert,
``The complete one-loop dilatation operator of N = 4 super Yang-Mills
theory,''
Nucl.\ Phys.\ B {\bf 676} (2004) 3
[arXiv:hep-th/0307015].
$\bullet$
N.~Beisert and M.~Staudacher,
``The N = 4 SYM integrable super spin chain,''
Nucl.\ Phys.\ B {\bf 670} (2003) 439
[arXiv:hep-th/0307042].
$\bullet$
N.~Beisert,
``The su(2$|$3) dynamic spin chain,''
Nucl.\ Phys.\ B {\bf 682} (2004) 487
[arXiv:hep-th/0310252].
$\bullet$
A.~A.~Tseytlin,
``Spinning strings and AdS/CFT duality,''
arXiv:hep-th/0311139.
%%CITATION = HEP-TH/0311139;%%
$\bullet$
V.~A.~Kazakov, A.~Marshakov, J.~A.~Minahan and K.~Zarembo,
``Classical / quantum integrability in AdS/CFT,''
JHEP {\bf 0405} (2004) 024
[arXiv:hep-th/0402207].
$\bullet$
N.~Beisert, V.~Dippel and M.~Staudacher,
``A novel long range spin chain and planar N = 4 super Yang-Mills,''
JHEP {\bf 0407} (2004) 075
[arXiv:hep-th/0405001].
$\bullet$
G.~Arutyunov, S.~Frolov and M.~Staudacher,
``Bethe ansatz for quantum strings,''
JHEP {\bf 0410} (2004) 016
[arXiv:hep-th/0406256].
$\bullet$
M.~Staudacher,
``The factorized S-matrix of CFT/AdS,''
JHEP {\bf 0505} (2005) 054
[arXiv:hep-th/0412188].
$\bullet$
N.~Beisert and M.~Staudacher,
``Long-range PSU(2,2$|$4) Bethe ansaetze for gauge theory and strings,''
Nucl.\ Phys.\ B {\bf 727} (2005) 1
[arXiv:hep-th/0504190].
$\bullet$
S.~Sch\"afer-Nameki, M.~Zamaklar and K.~Zarembo,
``Quantum corrections to spinning strings in AdS(5) x S**5 and Bethe
ansatz:
A comparative study,''
JHEP {\bf 0509} (2005) 051
[arXiv:hep-th/0507189].
$\bullet$
S.~Frolov, J.~Plefka and M.~Zamaklar,
``The AdS(5) x S**5 superstring in light-cone gauge and its Bethe
equations,''
J.\ Phys.\ A  {\bf 39} (2006) 13037
[arXiv:hep-th/0603008].
%%CITATION = JPAGB,A39,13037;%%
$\bullet$
R.~Hernandez and E.~Lopez,
``Quantum corrections to the string Bethe ansatz,''
JHEP {\bf 0607} (2006) 004
[arXiv:hep-th/0603204].
$\bullet$
L.~Freyhult and C.~Kristjansen,
``A universality test of the quantum string Bethe ansatz,''
Phys.\ Lett.\ B {\bf 638} (2006) 258
[arXiv:hep-th/0604069].
$\bullet$
D.~M.~Hofman and J.~M.~Maldacena,
``Giant magnons,''
J.\ Phys.\ A  {\bf 39}, 13095 (2006)
[arXiv:hep-th/0604135].
%%CITATION = JPAGB,A39,13095;%%
$\bullet$
T.~Klose, T.~McLoughlin, R.~Roiban and K.~Zarembo,
``Worldsheet scattering in AdS(5) x S**5,''
JHEP {\bf 0703} (2007) 094
[arXiv:hep-th/0611169].

\bibitem{AF}
G.~Arutyunov and S.~Frolov,
``On AdS(5) x S**5 string S-matrix,''
Phys.\ Lett.\ B {\bf 639} (2006) 378
[arXiv:hep-th/0604043].

\bibitem{BES}
N.~Beisert, R.~Hernandez and E.~Lopez,
``A crossing-symmetric phase for AdS(5) x S**5 strings,''
JHEP {\bf 0611} (2006) 070
[arXiv:hep-th/0609044].
$\bullet$
N.~Beisert, B.~Eden and M.~Staudacher,
``Transcendentality and crossing,''
J.\ Stat.\ Mech.\ {\bf 0701}, P021 (2007)
[arXiv:hep-th/0610251].
%%CITATION = JSTAT,0701,P021;%%

\bibitem{BBKS}
Z.~Bern, M.~Czakon, L.~J.~Dixon, D.~A.~Kosower and V.~A.~Smirnov,
``The Four-Loop Planar Amplitude and Cusp Anomalous Dimension in
Maximally Supersymmetric Yang-Mills Theory,''
Phys.\ Rev.\  D {\bf 75}, 085010 (2007)
[arXiv:hep-th/0610248].
%%CITATION = PHRVA,D75,085010;%%
$\bullet$
F.~Cachazo, M.~Spradlin and A.~Volovich,
``Four-Loop Cusp Anomalous Dimension From Obstructions,''
Phys.\ Rev.\  D {\bf 75}, 105011 (2007)
[arXiv:hep-th/0612309].
%%CITATION = PHRVA,D75,105011;%%
$\bullet$
M.~K.~Benna, S.~Benvenuti, I.~R.~Klebanov and A.~Scardicchio,
``A test of the AdS/CFT correspondence using high-spin operators,''
Phys.\ Rev.\ Lett.\ {\bf 98}, 131603 (2007)
[arXiv:hep-th/0611135].
%%CITATION = PRLTA,98,131603;%%
$\bullet$
L.~F.~Alday, G.~Arutyunov, M.~K.~Benna, B.~Eden and I.~R.~Klebanov,
``On the strong coupling scaling dimension of high spin operators,''
JHEP {\bf 0704}, 082 (2007)
[arXiv:hep-th/0702028].
%%CITATION = JHEPA,0704,082;%%
$\bullet$
I.~Kostov, D.~Serban and D.~Volin,
``Strong coupling limit of Bethe ansatz equations,''
arXiv:hep-th/0703031.
%%CITATION = HEP-TH/0703031;%%
$\bullet$
M.~Beccaria, G.~F.~De Angelis and V.~Forini,
``The scaling function at strong coupling from the quantum string Bethe 
equations,''
JHEP {\bf 0704}, 066 (2007)
[arXiv:hep-th/0703131].
%%CITATION = JHEPA,0704,066;%%
$\bullet$
P.~Y.~Casteill and C.~Kristjansen,
``The Strong Coupling Limit of the Scaling Function from the Quantum
String Bethe Ansatz,''
arXiv:0705.0890 [hep-th].
%%CITATION = ARXIV:0705.0890;%%
$\bullet$
D.~Fioravanti and M.~Rossi,
``On the commuting charges for the highest dimension SU(2) operator in
planar ${\cal N}=4$ SYM,''
arXiv:0706.3936 [hep-th].
%%CITATION = ARXIV:0706.3936;%%

\bibitem{Thomas}
V.~Giangreco Marotta Puletti, T.~Klose and O.~Ohlsson Sax,
``Factorized world-sheet scattering in near-flat $AdS_5$ x $S^5$,''
arXiv:0707.2082 [hep-th].
%%CITATION = ARXIV:0707.2082;%%

\bibitem{MS}
J.~Maldacena and I.~Swanson,
``Connecting giant magnons to the pp-wave: An interpolating limit of
$AdS_5 \times S^5$,''
Phys.\ Rev.\  D {\bf 76}, 026002 (2007)
[arXiv:hep-th/0612079].
%%CITATION = PHRVA,D76,026002;%%
$\bullet$
T.~Klose, T.~McLoughlin, J.~A.~Minahan and K.~Zarembo,
``World-sheet scattering in AdS(5) x S**5 at two loops,''
arXiv:0704.3891 [hep-th].
%%CITATION = ARXIV:0704.3891;%%

\bibitem{symm}
G.~Mandal, N.~V.~Suryanarayana and S.~R.~Wadia,
``Aspects of semiclassical strings in AdS(5),''
Phys.\ Lett.\  B {\bf 543}, 81 (2002)
[arXiv:hep-th/0206103].
%%CITATION = PHLTA,B543,81;%%
$\bullet$
I.~Bena, J.~Polchinski and R.~Roiban,
``Hidden symmetries of the AdS(5) x S**5 superstring,''
Phys. \ Rev. \ D {\bf 69} (2004) 046002
[arXiv:hep-th/0305116].
$\bullet$
L.~Dolan, C.~R.~Nappi and E.~Witten,
``Yangian symmetry in D = 4 superconformal Yang-Mills theory,''
arXiv:hep-th/0401243.
$\bullet$
M.~Hatsuda and K.~Yoshida,
``Classical integrability and super Yangian of superstring on AdS(5) x
S**5,''
Adv.\ Theor.\ Math.\ Phys.\  {\bf 9}, 703 (2005)
[arXiv:hep-th/0407044].
%%CITATION = 00203,9,703;%%
$\bullet$
A.~Das, J.~Maharana, A.~Melikyan and M.~Sato,
``The algebra of transition matrices for the AdS(5) x S**5
superstring,''
JHEP {\bf 0412}, 055 (2004)
[arXiv:hep-th/0411200].
%%CITATION = JHEPA,0412,055;%%
$\bullet$
A.~Das, A.~Melikyan and M.~Sato,
``The algebra of flat currents for the string on AdS(5) x S**5 in the
light-cone gauge,''
JHEP {\bf 0511}, 015 (2005)
[arXiv:hep-th/0508183].
%%CITATION = JHEPA,0511,015;%%
$\bullet$
A.~Agarwal,
``Comments on higher loop integrability in the su(1$|$1) sector of N = 4
SYM: Lessons from the su(2) sector,''
arXiv:hep-th/0506095.
%%CITATION = HEP-TH/0506095;%%
$\bullet$
B.~I.~Zwiebel,
``Yangian symmetry at two-loops for the su(2$|$1) sector of N = 4 SYM,''
J.\ Phys.\ A  {\bf 40}, 1141 (2007)
[arXiv:hep-th/0610283].
%%CITATION = JPAGB,A40,1141;%%
$\bullet$
C.~Gomez and R.~Hernandez,
``Quantum deformed magnon kinematics,''
JHEP {\bf 0703}, 108 (2007)
[arXiv:hep-th/0701200].
$\bullet$
C.~A.~S.~Young,
``q-Deformed Supersymmetry and Dynamic Magnon Representations,''
J.\ Phys.\ A  {\bf 40}, 9165 (2007)
[arXiv:0704.2069 [hep-th]].
%%CITATION = JPAGB,A40,9165;%%

\bibitem{FZ}
G.~Arutyunov, S.~Frolov and M.~Zamaklar,
``The Zamolodchikov-Faddeev algebra for AdS(5) x S**5 superstring,''
JHEP {\bf 0704}, 002 (2007)
[arXiv:hep-th/0612229].
%%CITATION = JHEPA,0704,002;%%

\bibitem{BZ}
N.~Beisert and B.~I.~Zwiebel,
``On Symmetry Enhancement in the psu(1,1$|$2) Sector of N=4 SYM,''
arXiv:0707.1031 [hep-th].
%%CITATION = ARXIV:0707.1031;%%

\bibitem{BeisYang}
N.~Beisert,
``The S-Matrix of AdS/CFT and Yangian Symmetry,''
PoS {\bf SOLVAY}, 002 (2006)
arXiv:0704.0400 [nlin.SI].
%%CITATION = POSCI,SOLVAY,002;%%

\bibitem{Dr1}
V.~G.~Drinfeld,
``Quantum groups,''
Proc. of the International Congress of Mathematicians, Berkeley,
1986, American Mathematical Society, 1987, 798.

\bibitem{Beinlin}
 N.~Beisert,
``The Analytic Bethe Ansatz for a Chain with Centrally Extended
su(2$|$2) Symmetry,''
J.\ Stat.\ Mech.\  {\bf 0701}, P017 (2007)
[arXiv:nlin.si/0610017].
%%CITATION = JSTAT,0701,P017;%%

\bibitem{Tor}
A.~Torrielli,
``Classical r-matrix of the su(2$|$2) SYM spin-chain,''
Phys. Rev. D {\bf 75} (2007) 105020
[arXiv:hep-th/0701281].

\bibitem{MorTor}
S.~Moriyama and A.~Torrielli,
``A Yangian Double for the AdS/CFT Classical r-matrix,''
JHEP {\bf 0706}, 083 (2007)
[arXiv:0706.0884 [hep-th]].
%%CITATION = JHEPA,0706,083;%%

\bibitem{Dr2}
V.~G.~Drinfeld,
``A new realization of Yangians and quantized affine algebras,''
Sov. Math. Dokl. {\bf 36} (1988) 212.
$\bullet$
I.~Heckenberger, F.~Spill, A.~Torrielli and H. Yamane, 
``Drinfeld second realization of the quantum affine superalgebras of
$D^{(1)}(2,1;x)$ via the Weyl groupoid,''
arXiv:0705.1071 [math.QA].

\bibitem{mst}
R.~Dijkgraaf and L.~Motl,
``Matrix string theory, contact terms, and superstring field theory,''
arXiv:hep-th/0309238.
%%CITATION = HEP-TH/0309238;%%
$\bullet$
S.~Moriyama,
``Comments on supersymmetry algebra and contact term in matrix string
theory,''
JHEP {\bf 0409}, 013 (2004)
[arXiv:hep-th/0405091].
%%CITATION = JHEPA,0409,013;%%
$\bullet$
I.~Kishimoto, S.~Moriyama and S.~Teraguchi,
``Twist field as three string interaction vertex in light cone string
field theory,''
Nucl.\ Phys.\  B {\bf 744}, 221 (2006)
[arXiv:hep-th/0603068].
%%CITATION = NUPHA,B744,221;%%
$\bullet$
I.~Kishimoto and S.~Moriyama,
``On LCSFT/MST correspondence,''
arXiv:hep-th/0611113.
%%CITATION = HEP-TH/0611113;%%

\bibitem{pp}
C.~Kristjansen, J.~Plefka, G.~W.~Semenoff and M.~Staudacher,
``A new double-scaling limit of N = 4 super Yang-Mills theory and
PP-wave strings,''
Nucl.\ Phys.\  B {\bf 643}, 3 (2002)
[arXiv:hep-th/0205033].
%%CITATION = NUPHA,B643,3;%%
$\bullet$
N.~R.~Constable, D.~Z.~Freedman, M.~Headrick, S.~Minwalla, L.~Motl,
A.~Postnikov and W.~Skiba,
``PP-wave string interactions from perturbative Yang-Mills theory,''
JHEP {\bf 0207}, 017 (2002)
[arXiv:hep-th/0205089].
%%CITATION = JHEPA,0207,017;%%
$\bullet$
D.~J.~Gross, A.~Mikhailov and R.~Roiban,
``A calculation of the plane wave string Hamiltonian from N = 4
super-Yang-Mills theory,''
JHEP {\bf 0305}, 025 (2003)
[arXiv:hep-th/0208231].
%%CITATION = JHEPA,0305,025;%%
$\bullet$
J.~Pearson, M.~Spradlin, D.~Vaman, H.~L.~Verlinde and A.~Volovich,
``Tracing the string: BMN correspondence at finite J**2/N,''
JHEP {\bf 0305}, 022 (2003)
[arXiv:hep-th/0210102].
%%CITATION = JHEPA,0305,022;%%
$\bullet$
J.~Gomis, S.~Moriyama and J.~w.~Park,
``SYM description of SFT Hamiltonian in a pp-wave background,''
Nucl.\ Phys.\  B {\bf 659}, 179 (2003)
[arXiv:hep-th/0210153].
%%CITATION = NUPHA,B659,179;%%
$\bullet$
J.~Gomis, S.~Moriyama and J.~w.~Park,
``SYM description of pp-wave string interactions: Singlet sector and
arbitrary impurities,''
Nucl.\ Phys.\  B {\bf 665}, 49 (2003)
[arXiv:hep-th/0301250].
%%CITATION = NUPHA,B665,49;%%
$\bullet$
P.~Gutjahr and A.~Pankiewicz,
``New aspects of the BMN correspondence beyond the planar limit,''
Nucl.\ Phys.\  B {\bf 704}, 583 (2005)
[arXiv:hep-th/0407098].
%%CITATION = NUPHA,B704,583;%%
$\bullet$
G.~Grignani, M.~Orselli, B.~Ramadanovic, G.~W.~Semenoff and D.~Young,
``Divergence cancellation and loop corrections in string field theory on
a plane wave background,''
JHEP {\bf 0512}, 017 (2005)
[arXiv:hep-th/0508126].
%%CITATION = JHEPA,0512,017;%%

\bibitem{susy}
P.~Di Vecchia, J.~L.~Petersen, M.~Petrini, R.~Russo and A.~Tanzini,
``The 3-string vertex and the AdS/CFT duality in the pp-wave limit,''
Class.\ Quant.\ Grav.\  {\bf 21}, 2221 (2004)
[arXiv:hep-th/0304025].
%%CITATION = CQGRD,21,2221;%%
$\bullet$
S.~Dobashi and T.~Yoneya,
``Resolving the holography in the plane-wave limit of AdS/CFT
correspondence,''
Nucl.\ Phys.\  B {\bf 711}, 3 (2005)
[arXiv:hep-th/0406225].
%%CITATION = NUPHA,B711,3;%%
$\bullet$
S.~Lee and R.~Russo,
``Holographic cubic vertex in the pp-wave,''
Nucl.\ Phys.\  B {\bf 705}, 296 (2005)
[arXiv:hep-th/0409261].
%%CITATION = NUPHA,B705,296;%%

\bibitem{SVPS}
M.~Spradlin and A.~Volovich,
``Superstring interactions in a pp-wave background,''
Phys.\ Rev.\  D {\bf 66}, 086004 (2002)
[arXiv:hep-th/0204146].
%%CITATION = PHRVA,D66,086004;%%
$\bullet$
M.~Spradlin and A.~Volovich,
``Superstring interactions in a pp-wave background. II,''
JHEP {\bf 0301}, 036 (2003)
[arXiv:hep-th/0206073].
%%CITATION = JHEPA,0301,036;%%
$\bullet$
A.~Pankiewicz and B.~J.~Stefanski,
``pp-wave light-cone superstring field theory,''
Nucl.\ Phys.\  B {\bf 657}, 79 (2003)
[arXiv:hep-th/0210246].
%%CITATION = NUPHA,B657,79;%%

\bibitem{MM}
M.~J.~Martins and C.~S.~Melo,
``The Bethe ansatz approach for factorizable centrally extended
S-matrices,''
arXiv:hep-th/0703086.
%%CITATION = HEP-TH/0703086;%%
\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\allowdisplaybreaks
\numberwithin{equation}{section}

\newcommand{\cS}{{\mathcal S}}
\newcommand{\cR}{{\mathcal R}}
\newcommand{\cC}{{\mathcal C}}
\newcommand{\cJ}{{\mathcal J}}
\newcommand{\fL}{{\mathfrak L}}
\newcommand{\fR}{{\mathfrak R}}
\newcommand{\fQ}{{\mathfrak Q}}
\newcommand{\fS}{{\mathfrak S}}
\newcommand{\fC}{{\mathfrak C}}
\newcommand{\fI}{{\mathfrak I}}
\newcommand{\fU}{{\mathfrak U}}
\newcommand{\fP}{{\mathfrak P}}
\newcommand{\fK}{{\mathfrak K}}
\newcommand{\alg}[1]{\mathfrak{#1}}

\begin{document}

\baselineskip=16pt plus 0.2pt minus 0.1pt

\begin{titlepage}
\title{
\hfill\parbox{4cm}
{\normalsize MIT-CTP-3853}\\
\vspace{1cm}
{\Large\bf A Secret Symmetry of the AdS/CFT S-matrix}
}
\author{
{\sc Takuya Matsumoto}$^{1}$\,{}\thanks
{{\tt m05044c@math.nagoya-u.ac.jp}}\\[6pt]
{\sc Sanefumi Moriyama}${}^{1,2}$\,{}\thanks
{{\tt moriyama@math.nagoya-u.ac.jp}}\\[6pt]
{\sc Alessandro Torrielli}${}^{2}$\,{}\thanks
{{\tt torriell@mit.edu}}\\[12pt]
${}^{1}${\it Graduate School of Mathematics, Nagoya University,}\\
{\it Nagoya 464-8602, Japan}\\[6pt]
${}^{2}${\it Center for Theoretical Physics, 
Massachusetts Institute of Technology,}\\
{\it Cambridge, MA02139, USA}\\
}
\date{\normalsize August, 2007}
\maketitle
\thispagestyle{empty}

\begin{abstract}
\normalsize
We find a new quantum Yangian symmetry of the AdS/CFT S-matrix, which
complements the original $\alg{su(2|2)}$ symmetry to $\alg{gl(2|2)}$ and
does not have a Lie algebra analog.
Our finding is motivated by the Yangian double structure discovered at
the classical level.
\end{abstract}

\end{titlepage}

\section{Introduction}

One of the most important results in the study of the integrable spin
chain inspired by AdS/CFT \cite{MZ}\footnote{We refer the reader to the
reviews \cite{revs} and references therein.} is that the relevant
S-matrix \cite{BeiSmat} can be determined uniquely by the
$\alg{su(2|2)}$ Lie algebra symmetry of the problem, up to an overall
dressing factor.
This fact reduces the problem of the dynamics into a single function.
The dressing factor satisfies the crossing symmetry constraint
\cite{Jan} originating from the underlying Hopf algebra structure
\cite{Jan,hopf,PST}.
A remarkable solution to Janik's equation, reproducing the asymptotic
behavior in the weak and strong coupling region \cite{pert,AF}, was
recently proposed by \cite{BES} and passed highly non-trivial checks
\cite{BBKS}.

In order to gain a deeper understanding of the hidden algebra
responsible for such a structure\footnote{See \cite{Thomas} for a recent
test of integrability in the near-flat space limit \cite{MS}.}, it is
desirable to understand all the symmetries of the model
\cite{symm,FZ,BZ}.
In \cite{BeisYang} it was shown that the whole Lie algebra symmetry
$\alg{su(2|2)}$ is lifted to the infinite-dimensional Yangian symmetry
by generalizing the standard formula of Drinfeld's first realization of
Yangians \cite{Dr1}\footnote{Note that we can freely rescale $\hbar$ at
the present stage. We shall fix it later to a convenient value.}:
\begin{align}
\Delta\widehat\cJ^A=\widehat\cJ^A\otimes 1
+1\otimes\widehat\cJ^A
+\frac{i}{2}\hbar f^A_{BC}\cJ^B\otimes\cJ^C~,\quad
{\rm S}(\widehat\cJ^A)=-\widehat\cJ^A
+\frac{i}{4}\hbar f^A_{BC}f^{BC}_D\cJ^D~.
\label{DandS}
\end{align}
Since the Cartan matrix which is used to raise and lower the indices is
degenerate, one appealed to $\alg{sl(2)}$ automorphisms \cite{Beinlin}
which couple to the central charges.

In the study of the classical (near BMN) limit \cite{Tor} it was noticed
that one of the poles reveals the Casimir operator of the Lie algebra
$\alg{gl(2|2)}$, instead of the expected symmetry algebra
$\alg{su(2|2)}$.
Although this fact suggests that the model has a larger symmetry of
$\alg{gl(2|2)}$, clearly the additional generator $\fI$ which extends
$\alg{su(2|2)}$ to $\alg{gl(2|2)}$
\begin{align}
\fI|\phi^a\rangle=I|\phi^a\rangle~,\quad
\fI|\psi^\alpha\rangle=-I|\psi^\alpha\rangle~,
\label{fI}
\end{align}
is not a symmetry of the S-matrix, when equipped with a trivial
coproduct.

A possible resolution to this problem was subsequently found in
\cite{MorTor} by rewriting the classical r-matrix in the form of a
Yangian double (Drinfeld's second realization \cite{Dr2}).
There, one found it necessary to include an infinite family of
generators $\fI_n$ to be able to factorize the classical r-matrix.
Although the coefficient of the additional generators $\fI_n$ remains
non-trivial for the infinite-dimensional Yangian algebra, it actually
vanishes for the classical Lie algebra at level $n=0$.

Here we would like to continue the study of this additional symmetry.
Since it does not have a Lie algebraic analog, we will call it a secret
symmetry.
Starting from the first higher Yangian level, the coproducts are usually
non-trivial, and there is a chance for the new generators to be
symmetries.
In fact, we will find a quantum Yangian generator proportional to $\fI$,
equipped with a coproduct that makes it a symmetry of the S-matrix.
We shall exploit a hybrid version of the arguments adopted in
\cite{BeisYang} and \cite{MorTor}.
Namely, we apply the formulas \eqref{DandS} together with the fact that
the $\fI$ operator couples to the central charge $\fC$ \cite{MorTor}.
We find
\begin{align}
\Delta\widehat\fI=\widehat\fI\otimes 1+1\otimes\widehat\fI
+\frac{i}{2g}\bigl(\fQ^\alpha{}_a\fU^{-1}\otimes\fS^a{}_\alpha
+\fS^a{}_\alpha\fU^{+1}\otimes\fQ^\alpha{}_a\bigr)~,\quad
{\rm S}(\widehat\fI)=-\widehat\fI+\frac{2i}{g}\fC~,\label{SI}
\end{align}
where $\widehat\fI$ acts as $\fI$ in \eqref{fI}, with the eigenvalue $I$
replaced by $\widehat I$. In the main text we will show that the
coproduct $\Delta\widehat\fI$ is a symmetry of the S-matrix if we choose
\begin{align}
\widehat I=\frac{1}{4}(x^++x^--1/x^+-1/x^-)~.
\end{align}
The Hopf algebra structure associated to the
Lie algebra generators $\fQ^\alpha{}_a$, $\fS^a{}_\alpha$ and $\fC$ in
\eqref{SI} is as described in \cite{BeisYang}.
The braiding factor $\fU$ is a central element with eigenvalue
$U=\sqrt{x^+/x^-}$.
As from \eqref{DandS}, it is easy to check that the defining relation of
Hopf algebras $\mu\circ({\rm S}\otimes 1)\circ\Delta=\eta\circ\epsilon$
($\mu$ is the algebra multiplication) holds for $\widehat\fI$, with the
counit $\epsilon(\widehat\fI)=0$.

A natural question arises whether the generator $\widehat\fI$ we found
is a generator in Drinfeld's second realization, so that we can use it
directly to construct the universal R-matrix.
Taking the classical limit of this operator, which essentially consists
in replacing $x^\pm$ by the classical spectral parameter $x$
\cite{AF,Tor}, we find
\begin{align}
\widehat I\to\frac{1}{2}(x-1/x)~.
\end{align}
This coincides with the expression of the first level generator
$\fI_{n=1}$ found in \cite{MorTor}, where Drinfeld's second realization
was employed.
Therefore, it appears natural to expect that the operator $\widehat\fI$
is in the second realization.
However, as we will argue below, this is not likely to be the case.

Since the generators in Drinfeld's second realization are the ones
directly appearing in the universal form of the R-matrix, they have to
be consistent with the crossing equation found by Janik \cite{Jan}:
\begin{align}
({\cal{C}}^{-1}\otimes 1)\bigl[\cR(1/x_1^\pm,x_2^\pm)\bigr]^{st_1}
({\cal{C}}\otimes 1)\cR(x_1^\pm,x_2^\pm)=1~,
\label{Janiks}
\end{align}
where ${\cal{C}}$ is a (bosonic) charge conjugation matrix and $st_1$
denotes supertransposition in the first entry.
This equation originates from combining the knowledge of the crossing
symmetry transformation of the Lie algebra ($n=0$) generators with the
fundamental property $({\rm S}\otimes 1){\cal{R}}={\cal{R}}^{-1}$ of
quasi-triangular Hopf algebras.
On the other hand, if we assume that the universal R-matrix $\cal{R}$
admits some expansion in powers of the Yangian generators $\cJ_n$,
\eqref{Janiks} leads to the expectation that all $\cJ_n$'s have to
satisfy the antipode relation
\begin{align}
\label{antipoderel}
{\rm S}(\cJ_n(x^\pm))
={\cC}^{-1}\bigl[\cJ_n(1/x^\pm)\bigr]^{st}{\cC}~,
\end{align}
with one and the same charge conjugation $\cal{C}$.
Apparently, although the operator $\widehat\fI$ satisfies the relation
$\cC^{-1}[\widehat\fI(1/x^\pm)]^{st}\cC=-\widehat\fI(x^\pm)$, it does
not satisfy \eqref{antipoderel}.
Hence, it is difficult to expect this operator to be in the second 
realization\footnote{Nevertheless, we still find it a little bizarre
that \eqref{antipoderel} is not satisfied by this new symmetry. This
issue certainly deserves further investigation, on which we reserve to
come back in the future.}.

As a side remark, we notice that all the generators found in
\cite{MorTor} satisfy the classical antipode relation
\begin{align}
{\rm S}(\cJ_n(x))
={\cC}_0^{-1}\bigl[\cJ_n(1/x)\bigr]^{st}{\cC}_0~,
\label{clantipode}
\end{align}
where $\cC_0$ is the classical charge conjugation, and the classical
antipode is given by
\begin{equation}
{\rm S}(\cJ_n(x))=-\cJ_n(x)~. 
\end{equation} 
The explicit expression of $\cC_0$ is not necessary to show that the
classical antipode relation \eqref{clantipode} holds for the infinite
tower of generators, provided the relation is satisfied by the Lie
algebra generators at level $n=0$ \cite{Jan,PST,FZ}.

In the last part of the introduction, let us recall for convenience the
action of the supercharges\footnote{We follow the notation of
\cite{Beinlin}.
In particular, note that $g$ in \cite{Beinlin} is related to the one in
\cite{BeiSmat} by $g_{\cite{Beinlin}}=g_{\cite{BeiSmat}}/\sqrt{2}$.}
\begin{align}
\fQ^\alpha{}_a|\phi^b\rangle=a\delta^b_a|\psi^\alpha\rangle~,&\quad
\fQ^\alpha{}_a|\psi^\beta\rangle
=b\epsilon^{\alpha\beta}\epsilon_{ab}|\phi^b\rangle~,\nonumber\\
\fS^a{}_\alpha|\phi^b\rangle
=c\epsilon^{ab}\epsilon_{\alpha\beta}|\psi^\beta\rangle~,&\quad
\fS^a{}_\alpha|\psi^\beta\rangle=d\delta^\beta_\alpha|\phi^a\rangle~,
\end{align}
with $a$, $b$, $c$ and $d$ parameterized by
\begin{align}
a=\sqrt{g}\gamma~,\quad
b=\sqrt{g}\frac{\alpha}{\gamma}\biggl(1-\frac{x^+}{x^-}\biggr)~,\quad
c=\sqrt{g}\frac{i\gamma}{\alpha x^+}~,\quad
d=\sqrt{g}\frac{x^+}{i\gamma}\biggl(1-\frac{x^-}{x^+}\biggr)~,
\end{align}
and the R-matrix \cite{BeiSmat,Beinlin} that we will use in this paper:
\begin{align}
\cR_{12}|\phi^a_1\phi^b_2\rangle
&=\frac{1}{2}(A_{12}-B_{12})|\phi^a_1\phi^b_2\rangle
+\frac{1}{2}(A_{12}+B_{12})|\phi^b_1\phi^a_2\rangle
+\frac{1}{2}C_{12}\epsilon^{ab}\epsilon_{\alpha\beta}
|\psi^\alpha_1\psi^\beta_2\rangle~,\nonumber\\
\cR_{12}|\psi^\alpha_1\psi^\beta_2\rangle
&=-\frac{1}{2}(D_{12}-E_{12})|\psi^\alpha_1\psi^\beta_2\rangle
-\frac{1}{2}(D_{12}+E_{12})|\psi^\beta_1\psi^\alpha_2\rangle
-\frac{1}{2}F_{12}\epsilon^{\alpha\beta}\epsilon_{ab}
|\phi^a_1\phi^b_2\rangle~,\nonumber\\
\cR_{12}|\phi^a_1\psi^\beta_2\rangle
&=G_{12}|\phi^a_1\psi^\beta_2\rangle
+H_{12}|\psi^\beta_1\phi^a_2\rangle~,\nonumber\\
\cR_{12}|\psi^\alpha_1\phi^b_2\rangle
&=L_{12}|\psi^\alpha_1\phi^b_2\rangle
+K_{12}|\phi^b_1\psi^\alpha_2\rangle~.
\end{align}
The functions $A_{12},B_{12},\ldots$ are given by
\begin{align}
\label{func}
&A_{12}=\frac{x_2^+-x_1^-}{x_2^--x_1^+}~,\quad
B_{12}=\frac{x_2^+-x_1^-}{x_2^--x_1^+}
\biggl(1-2\frac{1-1/x_1^+x_2^-}{1-1/x_1^+x_2^+}
\frac{x_2^--x_1^-}{x_2^+-x_1^-}\biggr)~,\nonumber\\
&\qquad C_{12}=\frac{2\gamma_1\gamma_2U_2}{\alpha x_1^+x_2^+}
\frac{1}{1-1/x_1^+x_2^+}
\frac{x_2^--x_1^-}{x_2^--x_1^+}~,\nonumber\\
&D_{12}=-\frac{U_2}{U_1}~,\quad
E_{12}=-\frac{U_2}{U_1}
\biggl(1-2\frac{1-1/x_1^-x_2^+}{1-1/x_1^-x_2^-}
\frac{x_2^+-x_1^+}{x_2^--x_1^+}\biggr)~,\nonumber\\
&\qquad F_{12}=-\frac{2\alpha(x_1^+-x_1^-)(x_2^+-x_2^-)}
{\gamma_1\gamma_2U_1x_1^-x_2^-}
\frac{1}{1-1/x_1^-x_2^-}
\frac{x_2^+-x_1^+}{x_2^--x_1^+}~,\nonumber\\
&G_{12}=\frac{1}{U_1}\frac{x_2^+-x_1^+}{x_2^--x_1^+}~,\quad
H_{12}=\frac{\gamma_1U_2}{\gamma_2U_1}
\frac{x_2^+-x_2^-}{x_2^--x_1^+}~,\nonumber\\
&L_{12}=U_2\frac{x_2^--x_1^-}{x_2^--x_1^+}~,\quad
K_{12}=\frac{\gamma_2}{\gamma_1}
\frac{x_1^+-x_1^-}{x_2^--x_1^+}~,
\end{align}
where $U=\sqrt{x^+/x^-}$ and $x^\pm$ satisfy the relation
$x^++1/x^+-x^--1/x^-=i/g$.

\section{Secret symmetry}

In this section, we would like to show that the coproduct of the
additional operator
\begin{align}
\Delta\widehat\fI=\widehat\fI\otimes 1+1\otimes\widehat\fI
+\frac{i}{2g}\bigl(\fQ^\alpha{}_a\fU^{-1}\otimes\fS^a{}_\alpha
+\fS^a{}_\alpha\fU^{+1}\otimes\fQ^\alpha{}_a\bigr)~,
\end{align}
is an exact symmetry of the S-matrix:
\begin{align}
\bigl[\Delta\widehat\fI,\cS\bigr]=0~.
\label{comm}
\end{align}
This equation can be expressed in terms of the R-matrix $\cR=\Pi\cS$
($\Pi$ is the graded permutation operator) as
\begin{align}
\bigl[\widehat\fI\otimes 1+1\otimes\widehat\fI,\cR\bigr]
=\frac{i}{2g}\bigl[\{\fQ,\fS\}_\otimes\,\cR
+\cR\,\{\fQ,\fS\}_\otimes\bigr]~.
\label{symmetry}
\end{align}
Here the expression on the right-hand-side is a bookkeeping notation:
\begin{align}
&\{\fQ,\fS\}_\otimes\,\cR
=\bigl(\fQ^\alpha{}_a\otimes\fU^{+1}\fS^a{}_\alpha
+\fS^a{}_\alpha\otimes\fU^{-1}\fQ^\alpha{}_a\bigr)\cR~,\nonumber\\
&\cR\,\{\fQ,\fS\}_\otimes
=\cR\bigl(\fQ^\alpha{}_a\fU^{-1}\otimes\fS^a{}_\alpha
+\fS^a{}_\alpha\fU^{+1}\otimes\fQ^\alpha{}_a\bigr)~.
\end{align}
Note that the braiding factor of the first (second) line is in the
second (first) entry.

The first term on the right-hand-side of \eqref{symmetry} is given by
\begin{align}
&\{\fQ,\fS\}_\otimes\,\cR\,|\phi^a_1\phi^b_2\rangle
=-B_{12}(a_1c_2U_2+c_1a_2U_2^{-1})
\epsilon^{ab}\epsilon_{\alpha\beta}|\psi^\alpha_1\psi^\beta_2\rangle
\nonumber\\&\hspace{5cm}
+C_{12}(-b_1d_2U_2-d_1b_2U_2^{-1})
\bigl(|\phi^a_1\phi^b_2\rangle
-|\phi^b_1\phi^a_2\rangle\bigr)~,\nonumber\\
&\{\fQ,\fS\}_\otimes\,\cR\,|\psi^\alpha_1\psi^\beta_2\rangle
=E_{12}(-b_1d_2U_2-d_1b_2U_2^{-1})
\epsilon^{\alpha\beta}\epsilon_{ab}|\phi^a_1\phi^b_2\rangle
\nonumber\\&\hspace{5cm}
-F_{12}(a_1c_2U_2+c_1a_2U_2^{-1})
\bigl(|\psi^\alpha_1\psi^\beta_2\rangle
-|\psi^\beta_1\psi^\alpha_2\rangle\bigr)~,\nonumber\\
&\{\fQ,\fS\}_\otimes\,\cR\,|\phi^a_1\psi^\beta_2\rangle
=G_{12}(a_1d_2U_2+c_1b_2U_2^{-1})|\psi^\beta_1\phi^a_2\rangle
+H_{12}(-b_1c_2U_2-d_1a_2U_2^{-1})|\phi^a_1\psi^\beta_2\rangle~,
\nonumber\\
&\{\fQ,\fS\}_\otimes\,\cR\,|\psi^\alpha_1\phi^b_2\rangle
=L_{12}(-b_1c_2U_2-d_1a_2U_2^{-1})|\phi^b_1\psi^\alpha_2\rangle
+K_{12}(a_1d_2U_2+c_1b_2U_2^{-1})|\psi^\alpha_1\phi^b_2\rangle~,
\end{align}
while the second term is
\begin{align}
&\cR\,\{\fQ,\fS\}_\otimes\,|\phi^a_1\phi^b_2\rangle
=(a_1U_1^{-1}c_2+c_1U_1a_2)\bigl(E_{12}
\epsilon^{ab}\epsilon_{\alpha\beta}|\psi^\alpha_1\psi^\beta_2\rangle
-F_{12}\bigl(|\phi^a_1\phi^b_2\rangle
-|\phi^b_1\phi^a_2\rangle\bigr)\bigr)~,\nonumber\\
&\cR\,\{\fQ,\fS\}_\otimes\,|\psi^\alpha_1\psi^\beta_2\rangle
=(-b_1U_1^{-1}d_2-d_1U_1b_2)\bigl(-B_{12}
\epsilon^{\alpha\beta}\epsilon_{ab}|\phi^a_1\phi^b_2\rangle
+C_{12}\bigl(|\psi^\alpha_1\psi^\beta_2\rangle
-|\psi^\beta_1\psi^\alpha_2\rangle\bigr)\bigr)~,\nonumber\\
&\cR\,\{\fQ,\fS\}_\otimes\,|\phi^a_1\psi^\beta_2\rangle
=(a_1U_1^{-1}d_2+c_1U_1b_2)
\bigl(L_{12}|\psi^\beta_1\phi^\alpha_2\rangle
+K_{12}|\phi^a_1\psi^\beta_2\rangle\bigr)~,\nonumber\\
&\cR\,\{\fQ,\fS\}_\otimes\,|\psi^\alpha_1\phi^b_2\rangle
=(-b_1U_1^{-1}c_2-d_1U_1a_2)
\bigl(G_{12}|\phi^b_1\psi^\alpha_2\rangle
+H_{12}|\psi^\alpha_1\phi^b_2\rangle\bigr)~.
\end{align}
Using the relations
\begin{align}
&-B_{12}(a_1c_2U_2+c_1a_2U_2^{-1})+E_{12}(a_1U_1^{-1}c_2+c_1U_1a_2)
=2igC_{12}(\widehat I_1+\widehat I_2)~,\nonumber\\
&C_{12}(-b_1d_2U_2-d_1b_2U_2^{-1})-F_{12}(a_1U_1^{-1}c_2+c_1U_1a_2)
=0~,\nonumber\\
&E_{12}(-b_1d_2U_2-d_1b_2U_2^{-1})-B_{12}(-b_1U_1^{-1}d_2-d_1U_1b_2)
=2igF_{12}(\widehat I_1+\widehat I_2)~,\nonumber\\
&-F_{12}(a_1c_2U_2+c_1a_2U_2^{-1})+C_{12}(-b_1U_1^{-1}d_2-d_1U_1b_2)
=0~,\nonumber\\
&G_{12}(a_1d_2U_2+c_1b_2U_2^{-1})+L_{12}(a_1U_1^{-1}d_2+c_1U_1b_2)
=4igH_{12}(\widehat I_1-\widehat I_2)~,\nonumber\\
&H_{12}(-b_1c_2U_2-d_1a_2U_2^{-1})+K_{12}(a_1U_1^{-1}d_2+c_1U_1b_2)
=0~,\nonumber\\
&L_{12}(-b_1c_2U_2-d_1a_2U_2^{-1})+G_{12}(-b_1U_1^{-1}c_2-d_1U_1a_2)
=-4igK_{12}(\widehat I_1-\widehat I_2)~,\nonumber\\
&K_{12}(a_1d_2U_2+c_1b_2U_2^{-1})+H_{12}(-b_1U_1^{-1}c_2-d_1U_1a_2)
=0~,
\label{computation}
\end{align}
if we define $\widehat I_{1(2)}$ as
\begin{align}
\widehat I=\frac{1}{4}(x^++x^--1/x^+-1/x^-)~,
\end{align}
we find
\begin{align}
&\bigl[\{\fQ,\fS\}_\otimes\,\cR+\cR\,\{\fQ,\fS\}_\otimes\bigr]
|\phi^a_1\phi^b_2\rangle
=2igC_{12}(\widehat I_1+\widehat I_2)
\epsilon^{ab}\epsilon_{\alpha\beta}
|\psi^\alpha_1\psi^\beta_2\rangle~,\nonumber\\
&\bigl[\{\fQ,\fS\}_\otimes\,\cR+\cR\,\{\fQ,\fS\}_\otimes\bigr]
|\psi^\alpha_1\psi^\beta_2\rangle
=2igF_{12}(\widehat I_1+\widehat I_2)
\epsilon^{\alpha\beta}\epsilon_{ab}
|\phi^a_1\phi^b_2\rangle~,\nonumber\\
&\bigl[\{\fQ,\fS\}_\otimes\,\cR+\cR\,\{\fQ,\fS\}_\otimes\bigr]
|\phi^a_1\psi^\beta_2\rangle
=4igH_{12}(\widehat I_1-\widehat I_2)
|\psi^\beta_1\phi^a_2\rangle~,\nonumber\\
&\bigl[\{\fQ,\fS\}_\otimes\,\cR+\cR\,\{\fQ,\fS\}_\otimes\bigr]
|\psi^\alpha_1\phi^b_2\rangle
=-4igK_{12}(\widehat I_1-\widehat I_2)
|\phi^b_1\psi^\alpha_2\rangle~.
\label{result}
\end{align}

Most of the computations in \eqref{computation} are tedious but
straightforward, except the first and the third line with $B_{12}$ and
$E_{12}$, which require a more elaborated treatment.
The easiest way to compute them is to separately deal with the `$1$'
term and the `$-2$' term in the parenthesis of $B_{12}$ and $E_{12}$ in
\eqref{func}.
It is not difficult to find that the `$-2$' term is proportional to
$C_{12}$ and $F_{12}$ respectively.
For the `$1$' term, the following identity is useful
\begin{align}
x_2^+x_2^--x_1^+x_1^-=2\frac{x_2^--x_1^-}{1-1/x_1^+x_2^+}
(\widehat I_1+\widehat I_2)~,
\end{align}
which can be obtained by multiplying the identity
\begin{align}
x_2^+x_2^--x_1^+x_1^-=\frac{1}{2}
\bigl((x_2^++x_1^+)(x_2^--x_1^-)+(x_2^-+x_1^-)(x_2^+-x_1^+)\bigr)~,
\end{align}
by the quantity
$(1-1/x_1^+x_2^+)/(x_2^--x_1^-)=(1-1/x_1^-x_2^-)/(x_2^+-x_1^+)$.

Comparing \eqref{result} with
\begin{align}
&\bigl[\widehat\fI\otimes 1+1\otimes\widehat\fI,\cR\bigr]
|\phi^a_1\phi^b_2\rangle
=2(-\widehat I_1-\widehat I_2)(C_{12}/2)
\epsilon^{ab}\epsilon_{\alpha\beta}
|\psi^\alpha_1\psi^\beta_2\rangle~,\nonumber\\
&\bigl[\widehat\fI\otimes 1+1\otimes\widehat\fI,\cR\bigr]
|\psi^\alpha_1\psi^\beta_2\rangle
=2(\widehat I_1+\widehat I_2)(-F_{12}/2)
\epsilon^{\alpha\beta}\epsilon_{ab}
|\phi^a_1\phi^b_2\rangle~,\nonumber\\
&\bigl[\widehat\fI\otimes 1+1\otimes\widehat\fI,\cR\bigr]
|\phi^a_1\psi^\beta_2\rangle
=2(-\widehat I_1+\widehat I_2)H_{12}
|\psi^\beta_1\phi^a_2\rangle~,\nonumber\\
&\bigl[\widehat\fI\otimes 1+1\otimes\widehat\fI,\cR\bigr]
|\psi^\alpha_1\phi^b_2\rangle
=2(\widehat I_1-\widehat I_2)K_{12}
|\phi^b_1\psi^\alpha_2\rangle~,
\end{align}
we eventually find that \eqref{symmetry} holds.

\section{Conclusion}

We have found a new quantum Yangian symmetry of the AdS/CFT S-matrix,
which complements the original $\alg{su(2|2)}$ symmetry to
$\alg{gl(2|2)}$ and does not have a Lie algebra analog.
Using this novel symmetry we can generate several new ones.
For example, we can compute the commutators
$[\Delta\widehat\fI,\Delta\fQ^\alpha{}_{b}]$ and
$[\Delta\widehat\fI,\Delta\fS^a{}_{\beta}]$.
Taking the linear combination with the known invariances of the S-matrix
$\Delta\widehat\fQ^\alpha{}_b$ and $\Delta\widehat\fS^a{}_\beta$
discovered in \cite{BeisYang}, we find new symmetries\footnote{As a
check, we have explicitly verified that one of them, namely
$\Delta\fQ^\alpha{}_{b,1}$, is an exact symmetry of the S-matrix:
$[\Delta\fQ^\alpha{}_{b,1},\cS]=0$.}:
($\hbar=1/(2g)$)
\begin{align}
&\Delta\fQ^\alpha{}_{b,1}
=\fQ^\alpha{}_{b,1}\otimes 1
+\fU^{+1}\otimes\fQ^\alpha{}_{b,1}
+i\hbar
\bigl(\fL^\alpha{}_\gamma\fU^{+1}\otimes\fQ^\gamma{}_b
+\fR^c{}_b\fU^{+1}\otimes\fQ^\alpha{}_c
+\fC\fU^{+1}\otimes\fQ^\alpha{}_b\nonumber\\
&\hspace{7cm}
-\fQ^\gamma{}_b\otimes\fL^\alpha{}_\gamma
-\fQ^\alpha{}_c\otimes\fR^c{}_b
-\fQ^\alpha{}_b\otimes\fC\bigr)~,
\nonumber\\
&\Delta\fQ^\alpha{}_{b,-1}
=\fQ^\alpha{}_{b,-1}\otimes 1
+\fU^{+1}\otimes\fQ^\alpha{}_{b,-1}
+i\hbar\bigl(
-\epsilon^{\alpha\gamma}\epsilon_{bd}\fP\fU^{-1}\otimes\fS^d{}_\gamma
+\epsilon^{\alpha\gamma}\epsilon_{bd}\fS^d{}_\gamma\fU^{+2}\otimes\fP
\bigr)~,
\nonumber\\
&\Delta\fS^a{}_{\beta,1}
=\fS^a{}_{\beta,1}\otimes 1
+\fU^{-1}\otimes\fS^a{}_{\beta,1}
+i\hbar
\bigl(-\fR^a{}_c\fU^{-1}\otimes\fS^c{}_\beta
-\fL^\gamma{}_\beta\fU^{-1}\otimes\fS^a{}_\gamma
-\fC\fU^{-1}\otimes\fS^a{}_\beta\nonumber\\
&\hspace{7cm}+\fS^c{}_\beta\otimes\fR^a{}_c
+\fS^a{}_\gamma\otimes\fL^\gamma{}_\beta
+\fS^a{}_\beta\otimes\fC\bigr)~,
\nonumber\\
&\Delta\fS^a{}_{\beta,-1}
=\fS^a{}_{\beta,-1}\otimes 1
+\fU^{-1}\otimes\fS^a{}_{\beta,-1}
+i\hbar\bigl(
\epsilon^{ac}\epsilon_{\beta\delta}\fK\fU^{+1}\otimes\fQ^\delta{}_c
-\epsilon^{ac}\epsilon_{\beta\delta}\fQ^\delta{}_c\fU^{-2}\otimes\fK
\bigr)~,
\end{align}
where $\fQ^\alpha{}_{b,\pm 1}$ and $\fS^a{}_{\beta,\pm 1}$ are defined
as
\begin{align}
\fQ^\alpha{}_{b,1}=\fQ^\alpha{}_b
\bigl(u\Pi_{\rm b}+v\Pi_{\rm f}\bigr)~,&\quad
\fQ^\alpha{}_{b,-1}=\fQ^\alpha{}_b
\bigl(v\Pi_{\rm b}+u\Pi_{\rm f}\bigr)~,\nonumber\\
\fS^a{}_{\beta,1}=\fS^a{}_\beta
\bigl(v\Pi_{\rm b}+u\Pi_{\rm f}\bigr)~,&\quad
\fS^a{}_{\beta,-1}=\fS^a{}_\beta
\bigl(u\Pi_{\rm b}+v\Pi_{\rm f}\bigr)~,
\end{align}
with $\Pi_{\rm b}$ ($\Pi_{\rm f}$) being the projector to bosons
(fermions) and
\begin{align}
u=\frac{1}{2}(x^++x^-)~,\quad v=\frac{1}{2}(1/x^++1/x^-)~.
\end{align}
The antipodes are given by 
${\rm S}(\fQ^\alpha{}_{b,\pm 1})=-\fU^{-1}\fQ^\alpha{}_{b,\pm 1}$ and
${\rm S}(\fS^a{}_{\beta,\pm 1})=-\fU^{+1}\fS^a{}_{\beta,\pm 1}$.
The names we have chosen for these operators are inspired by their
classical limit, and by the fact that they fulfill the crossing symmetry
property \eqref{antipoderel} at the quantum level.
Further commutators such as
$\{\Delta\fQ^\alpha{}_{b,1},\Delta\fS^a{}_\beta\}$ or
$\{\Delta\fQ^\alpha{}_b,\Delta\fS^a{}_{\beta,1}\}$ actually generate
the coproducts $\Delta\widehat\fR^a{}_b$ and
$\Delta\widehat\fL^\alpha{}_\beta$ of \cite{BeisYang}, which satisfy the
crossing symmetry relation at the quantum level, and coincide with the
generators $\Delta\fR^a{}_{b,1}$ and $\Delta\fL^\alpha{}_{\beta,1}$
found in \cite{MorTor} in the classical limit.

Let us list several further directions related to the results we have
presented here.

It is always crucial to study the symmetries when investigating the
correspondence between two theories.
The primary question which comes to mind is of course to see whether and
how the new one we have found in this paper will appear in the string
worldsheet theory.

An interesting application might be related to the string interaction
vertices.
In discussing the relation \cite{mst} between matrix string theory and
light-cone string field theory on the flat space, the supersymmetry
algebra on both sides plays an important role.
When it comes to the AdS/CFT correspondence on the pp-wave background at
the interacting level \cite{pp}, the story becomes more complicated.
It was questioned \cite{susy} whether supersymmetry alone can determine
the whole string interaction vertex \cite{SVPS}.
We would like to see whether our new symmetry can give further
restrictions in its construction.

Another application is related to the Hubbard model.
It was noticed \cite{Beinlin,MM} that the AdS/CFT S-matrix is equivalent
to Shastry's S-matrix.
Understanding how this new hidden symmetry appears in the Hubbard model
could help understanding more deeply the observed equivalence.

Finally, another interesting direction is to explore whether this
symmetry could be embedded in the full $\alg{psu(2,2|4)}$ algebra before
breaking it to $(\alg{su(2|2)})^2$, when the ground state is identified.
We reserve to return on these issues in future investigations.

\section*{Acknowledgments}
We would like to warmly thank G.~Arutyunov, H.~Awata, N.~Beisert,
S.~Benvenuti, H.Y.~Chen, S.~Frolov, J.~Gomis, H.~Kanno, Y.~Kazama,
T.~Klose, S.~Lee, X.~Liu, S.~Matsuura, T.~McLoughlin, J.~Minahan,
F.~Ravanini, S.~Sch\"afer-Nameki, K.~Yoshida and C.~Young for
interesting discussions.
A.T.\ especially thanks F.~Spill for many discussions and
for sharing his Diploma thesis.
S.M.\ is grateful for warm hospitality to the organizers of the
``Strings 007'' conference at Benasque, where part of this work was
done.
This work is supported in part by funds provided by the U.S. 
Department of Energy (D.O.E.) under cooperative research agreement
DE-FG02-05ER41360.
The work of S.M. is supported partly by Nishina Memorial Foundation,
Inamori Foundation and Grant-in-Aid for Young Scientists (\#18740143)
from the Japan Ministry of Education, Culture, Sports, Science and
Technology.
A.T. thanks Istituto Nazionale di Fisica Nucleare (I.N.F.N.) for
supporting him through a ``Bruno Rossi'' postdoctoral fellowship.

\end{document}